OpenStudy (anonymous):

need help............. Reduce all fractional answers to lowest terms. A line that passes midway between and is parallel to both x - y = 6 and x - y = -6.

OpenStudy (***[isuru]***):

hi, let's take the first one.. x - y = 6 y = x - 6 and the second one.. x - y = -6 y = x + 6 hence the line question asks ask to draw should be parallel to these 2 lines the slope should not be changed. which means.. If the line we have to draw have the shape.. y = mx + c where m is the gradient or slope and the c is y intercept m should be 1 cause the "m" (slope) of the given 2 is also 1 then we can short the line's equation y = x + c the next condition is .. that the line has to pass midway in between given 2 the first line goes through the -6 when x = 0 the second goes through the 6 when x = 0 the mid point between -6 and 6 is 0 then the line we have to draw should pass 0 when x = 0 which means it's y intercept(c) should be 0 then the function of the line is... y= x hope this will help ya!!

OpenStudy (anonymous):

i did thankx :)

OpenStudy (***[isuru]***):

u r welcome!